OpenStudy (anonymous):

Simplify. Show all work please. 8-6+1/-4^2+7

OpenStudy (lilg132):

-1/16 + 7 + 8 -6 = -1/16 + 9 = 9 * 16 = 144 - 1 = 143/16 :D